OpenStudy (anonymous):

Use a system of equations to solve the following problem: Douglas High School has a total of 89 boys and girls who play sports. If the number of boys is 8 more than twice the number of girls, how many boys play sports at this high school?

OpenStudy (anonymous):

x = boys y = girls x+y=89 x=8+2y Substitution (8+2y)+y=89 8 + 3y = 89 3y = 81 y = 27 girls x=8 + 2(27) = 62 boys